OpenStudy (anonymous):

Medals! The quotient of (x^4 – 3x^2 + 4x – 3) and a polynomial is (x^2 + x – 3). What is the polynomial? x^4 – 2x^2 + 5x – 6 x^2 – x + 1 x^6 + x^5 – 6x^4 + x^3 + 10x^2 – 15x + 9 x^4 – 4x^2 + 3x

OpenStudy (anonymous):

\[ x^2+x-3 = \frac{x^4 – 3x^2 + 4x – 3}{p(x)} \]So: \[ p(x)[x^2+x-3] = x^4 – 3x^2 + 4x – 3 \]

OpenStudy (anonymous):

Also \[ p(x) = \frac{x^4 – 3x^2 + 4x – 3}{x^2 + x – 3} \]

OpenStudy (anonymous):

So you divide the dividend by the quotient to find the divisor?

OpenStudy (anonymous):

Hello? You there?